# A Comprehensive Survey of Scientific Large Language Models and Their Applications in Scientific Discovery

### v2

@@ Abstract
In many scientific fields, large language models (LLMs) have revolutionized the way text and other modalities of data (e.g., molecules and proteins) are handled, achieving superior performance in various applications and augmenting the scientific discovery process. Nevertheless, previous surveys on scientific LLMs often concentrate on one or two fields or a single modality. In this paper, we aim to provide a more holistic view of the research landscape by unveiling cross-field and cross-modal connections between scientific LLMs regarding their architectures and pre-training techniques. To this end, we comprehensively survey over 250 scientific LLMs, discuss their commonalities and differences, as well as summarize pre-training datasets and evaluation tasks for each field and modality. Moreover, we investigate how LLMs have been deployed to benefit scientific discovery. Resources related to this survey are available at https://github.com/yuzhimanhua/Awesome-Scientific-Language-Models. * Equal contribution
@@ 1 Introduction
Given the broad and profound impact of LLMs in various scientific fields across diverse modalities, it becomes necessary to comprehensively review related work in this direction. However, existing scientific LLM surveys typically focus on either one or two fields (e.g., biomedicine Wang et al. (2023a); He et al. (2024); Pei et al. (2024); Zhang et al. (2024d) and chemistry Xia et al. (2023); Pei et al. (2024); Zhang et al. (2024d)) or one modality (e.g., text Ho et al. (2024)) only. In fact, if we take a holistic view of the research landscape, we can observe similar and interrelated techniques used to develop LLMs for different fields and modalities.
@@
Figure 1 depicts three major types of scientific LLM pre-training strategies (i.e., Columns 1 to 3), for each of which we give 4 examples (i.e., Types a to d). In Column 1, following BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), existing studies use masked language modeling (MLM) to pre-train encoder language models. Here, the input can be naturally sequential (e.g., papers in each field, protein/DNA/RNA sequences in the FASTA format Lipman and Pearson (1985)) or artificially linearized (e.g., molecules in the SMILES format Weininger (1988), sequences of venue/author/reference nodes in citation graphs). In Column 2, inspired by GPT Brown et al. (2020) and LLaMA Touvron et al. (2023a), previous studies adopt next token prediction to pre-train (encoder-)decoder language models, some of which further adopt instruction tuning and preference optimization Ouyang et al. (2022). Other than plain text input (e.g., question-answer pairs from knowledge bases or exams), we see more ways to sequentialize complex scientific data, such as flattening table cells and using particle coordinates to describe crystals. Even for images, there are studies in both mathematics Gao et al. (2023) and biomedicine Li et al. (2023a) that exploit a vision encoder to project an image onto several visual tokens and prepend them to text tokens as linearized LLM input. In Column 3, following DPR Karpukhin et al. (2020) and CLIP Radford et al. (2021), two encoders are pre-trained to map relevant data pairs closer in the latent space via contrastive learning. When both modalities are sequential (e.g., text-text or text-protein), the model is built upon two LLM encoders. When we prefer to keep the non-sequential nature of one modality (e.g., molecular graphs Edwards et al. (2021), chest X-rays Zhang et al. (2022), and aerial views Yan et al. (2024)), the corresponding graph or image encoder can be employed. To summarize, a cross-field, cross-modal survey will more accurately draw the connections between different scientific LLMs, demonstrate their commonalities, and potentially guide their future designs.
@@
Contributions. In this paper, motivated by the discussions above, we systematically survey over 250 scientific LLMs encompassing various fields (e.g., general science, mathematics, physics, chemistry, materials science, biology, medicine, and geoscience), modalities (e.g., language, graph, vision, table, molecule, protein, genome, and climate time series), and sizes (from $\sim$100M to $\sim$100B parameters). For each field/modality, we investigate commonly adopted pre-training datasets, model architectures, and evaluation tasks of scientific LLMs. Following our motivation, when we discuss model architectures in detail, we link them back to Figure 1 to build cross-field cross-modal connections. Moreover, we provide a structured summary of these scientific LLMs in Table A1-Table A6 (Appendix A). Furthermore, for different fields, we introduce how LLMs have been deployed to benefit science by augmenting different aspects and stages of the scientific discovery process, such as hypothesis generation, theorem proving, experiment design, drug discovery, and weather forecasting.
@@ 2.1 Language
General scientific LLMs are usually evaluated on common NLP tasks, such as named entity recognition (NER), relation extraction (RE) Luan et al. (2018), question answering (QA) Wang et al. (2024e), and classification Cohan et al. (2019).
@@ 2.3 Applications in Scientific Discovery
Performant scientific LLMs can work alongside researchers throughout the entire scientific discovery process. Leaving field-specific applications for later sections, here we underscore LLMs’ general usefulness in brainstorming and evaluation: Lahav et al. (2022) integrate LLMs into a search engine for the discovery of scientific challenges and directions; Wang et al. (2023e), Yang et al. (2024d), Baek et al. (2024), and Gu and Krenn (2024) leverage LLMs to generate novel scientific ideas, directions, and hypotheses on the basis of prior literature and existing knowledge; Zhang et al. (2023h) rely on LLMs to find expert reviewers for each submission; Liu and Shah (2023), Liang et al. (2024c), and D’Arcy et al. (2024) explore the capacity of GPT-4 to provide useful feedback on research papers to facilitate automatic review generation; Liang et al. (2024b, a) also observe the increasing use of LLMs in writing scientific papers and conference peer reviews.
@@ 3.1 Language
The pre-training text corpora for math LLMs can be categorized into two classes: (1) multiple-choice QA, the representative datasets of which include MathQA Amini et al. (2019), Ape210K Zhao et al. (2020), and Math23K Wang et al. (2017); as well as (2) generative QA, the representative datasets of which include GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), and MetaMathQA Yu et al. (2024c).
Similarly to general science LLMs, the backbone model of pioneering math LLMs is BERT (Type 1.a, e.g., GenBERT Geva et al. (2020) and MathBERT Shen et al. (2021)). These models are mostly trained via MLM, with the only exception being BERT-TD Li et al. (2022c), where a contrastive loss is adopted. For GPT-based math LLMs (Type 2.a, e.g., GSM8K-GPT Cobbe et al. (2021b) and NaturalProver Welleck et al. (2022)), we find a diversity of pre-training tasks: supervised fine-tuning, next token prediction, and instruction tuning. The most recent math LLMs (Type 2.a, e.g., Rho-Math Lin et al. (2024b) and MAmmoTH2 Yue et al. (2024c)) are based on LLaMA and are trained to follow natural language instructions. However, when an enormous pre-training dataset (e.g., 55 billion tokens) is available, next token prediction is still favored as the mere pre-training task Azerbayev et al. (2024); Lin et al. (2024b) or the companion task Shao et al. (2024); Ying et al. (2024) to build base models.
QA and math world problems (MWP) have been the most common evaluation tasks. In addition, quantitative reasoning contains more difficult problems, as the model has to provide a complete and self-contained solution without relying on external tools Shao et al. (2024); Lin et al. (2024b). We see a dominance of use from GSM8K and MATH for QA, and from MathQA and Math23K for MWP. For quantitative reasoning, MMLU-STEM Hendrycks et al. (2020) and Big-Bench Hard Suzgun et al. (2023) are the most widely adopted.
@@ 3.2 Language + Vision
Geometry is one of the most important branches of mathematics, and it expresses the settings jointly in text and diagrams. As such, it is mandatory to involve the vision modality for geometric LLMs. The most commonly used pre-training datasets for geometric LLMs include Geometry3K Lu et al. (2021) and GeoQA Chen et al. (2021), both of which contain multiple-choice geometry problems.
@@
Geometric LLMs are evaluated through geometry problem solving, where the model is asked to select the correct answer given the diagram and its caption, the question, and answer options. Renowned evaluation datasets include Geometry3K Lu et al. (2021), GEOS Seo et al. (2015), and MathVista Lu et al. (2023b).
@@ 3.3 Table
The challenge in tables is similar to that in diagrams, namely to obtain linearized table representations. In most cases, tables are squeezed into linear text sequences as part of the context and are prepended with the question text as the model input. As one of the first works in this line of research, TAPAS Herzig et al. (2020) (Type 1.a) adopts the MLM objective to predict the masked token in textual and tabular context. The most recent developments Li et al. (2024b); Zhang et al. (2024f) resemble the design of TableLlama Zhang et al. (2024e) (Type 2.b), with LLaMA-2 as the backbone and instruction tuning as the pre-training task.
@@ 3.4 Applications in Scientific Discovery
Math LLMs have great potential to assist humans in offering potential solutions. For instance, AlphaGeometry Trinh et al. (2024) combines an LLM with a symbolic deduction engine, where the LLM generates useful constructs and the symbolic engine applies formal logic to find solutions. AlphaGeometry solves 25 out of 30 classical geometry problems adapted from the International Mathematical Olympiad. Sinha et al. (2024) extend AlphaGeometry by adding Wu’s method Chou (1988), further solving 27 out of 30, surpassing human gold medalists. FunSearch Romera-Paredes et al. (2024) integrates LLM with program search. One notable achievement of FunSearch is its ability to find a new solution to the cap set problem in combinatorial optimization. The solutions generated can be faster and more efficient than those devised by human experts. In Li et al. (2024a), LLMs iteratively propose and critique statistical models by leveraging in-context learning and chain-of-thought reasoning Wei et al. (2022b).
@@ 5.1 Language
LLM pre-training corpora in chemistry and materials science typically come from research papers and databases (e.g., Materials Project Jain et al. (2013)). Besides, recent works adopt domain-specific instruction tuning datasets (e.g., Mol-Instructions Fang et al. (2023) and SMolInstruct Yu et al. (2024a)) derived from PubChem Kim et al. (2019), MoleculeNet Wu et al. (2018), etc.
@@ 5.2 Language + Graph
In some scenarios, molecular graphs appear simultaneously with text information, thus existing works have explored how to encode both effectively. The first type of such models adopts a GNN as the graph encoder and an LLM as the text encoder. The two modalities are connected through contrastive learning Liu et al. (2023d) (Type 3.c). For example, Text2Mol Edwards et al. (2021) uses GCN Kipf and Welling (2016) and SciBERT to encode a molecule and its corresponding natural language description, respectively, for text-to-molecule retrieval. The second type of models utilizes an LLM to encode text and graphs simultaneously Zeng et al. (2022). Graphs can be either linearized to SMILES strings Edwards et al. (2022) (Type 2.c) or projected onto virtual tokens with graph encoders Zhao et al. (2023a); Liu et al. (2023e) (Type 2.d). For instance, 3D-MoLM Li et al. (2024c) uses a 3D molecular encoder to represent molecules as tokens, and feeds them together with instructions into LLaMA-2 for molecule-to-text retrieval and molecule captioning.
@@ 5.4 Molecule
Different from subsection 5.2, this subsection introduces models dealing with molecules without associated text information. That being said, comparable approaches inspired by LLMs are utilized to develop molecular language models Flam-Shepherd et al. (2022). To be specific, most studies adopt SMILES or SELFIES Krenn et al. (2020) strings as the sequential representation of molecules. Similar to the trend in the “Language” modality, pioneering molecular LLMs focus on representation learning with bidirectional Transformer encoders (Type 1.c, e.g., SMILES-BERT Wang et al. (2019) and MoLFormer Ross et al. (2022)). For instance, ChemBERTa Chithrananda et al. (2020) adopts the architecture and pre-training strategy similar to those of RoBERTa Liu et al. (2019). These models exhibit extraordinary abilities in molecular understanding tasks such as molecular property prediction (e.g., toxicity classification Wu et al. (2018) and atomization energy regression Ramakrishnan et al. (2014)) as well as virtual screening Riniker and Landrum (2013). Later works explore representing molecules in an autoregressive fashion (Type 2.c, e.g., BARTSmiles Chilingaryan et al. (2024) and ChemGPT Frey et al. (2023)). For instance, T5Chem Lu and Zhang (2022) adopts the T5 backbone and a sequence-to-sequence pre-training objective. These models are evaluated in generative tasks that include molecule generation Gaulton et al. (2017), reaction prediction, and retrosynthesis. Besides linearizing molecules, there are studies modifying the Transformer architecture to admit molecular graphs, such as MAT Maziarka et al. (2020) and R-MAT Maziarka et al. (2024).
@@ 6.1 Language
The watershed moment in the evolution biomedical LLMs is still the emergence of billion-parameter architectures and instruction tuning. Before that, a wide variety of moderate-sized backbones are explored, including both encoder-based (Type 1.a, e.g., BioBERT Lee et al. (2020), Bio-ELECTRA Ozyurt (2020), BioRoBERTa Lewis et al. (2020b), BioALBERT Naseem et al. (2022), and Clinical-Longformer Li et al. (2022a)) and (encoder-) decoder-based ones (Type 2.a, e.g., SciFive Phan et al. (2021), BioBART Yuan et al. (2022a), and BioGPT Luo et al. (2022)). Evaluation tasks for these models range from biomedical NER, RE, sentence similarity estimation, document classification, and QA (i.e., the BLURB benchmark Gu et al. (2021)) to natural language inference (NLI) Romanov and Shivade (2018) and entity linking Doğan et al. (2014). After the watershed, the trend becomes instruction-tuning billion-parameter LLMs (Type 2.a, e.g., Med-PaLM Singhal et al. (2023a), MedAlpaca Han et al. (2023), and BioMistral Labrak et al. (2024)). Accordingly, evaluation tasks now include single-round QA Jin et al. (2021); Pal et al. (2022) and multi-round dialogue Wang et al. (2023f). Meanwhile, there are studies proposing a Bi-Encoder architecture (Type 3.a, e.g., Jin et al. (2023c) and Xu et al. (2024)) that specifically targets biomedical retrieval tasks, the benchmarks of which are NFCorpus Boteva et al. (2016), TREC-COVID Voorhees et al. (2021), etc.

### v3

@@ Abstract
In many scientific fields, large language models (LLMs) have revolutionized the way text and other modalities of data (e.g., molecules and proteins) are handled, achieving superior performance in various applications and augmenting the scientific discovery process. Nevertheless, previous surveys on scientific LLMs often concentrate on one or two fields or a single modality. In this paper, we aim to provide a more holistic view of the research landscape by unveiling cross-field and cross-modal connections between scientific LLMs regarding their architectures and pre-training techniques. To this end, we comprehensively survey over 260 scientific LLMs, discuss their commonalities and differences, as well as summarize pre-training datasets and evaluation tasks for each field and modality. Moreover, we investigate how LLMs have been deployed to benefit scientific discovery. Resources related to this survey are available at https://github.com/yuzhimanhua/Awesome-Scientific-Language-Models. * Equal contribution
@@ 1 Introduction
Given the broad and profound impact of LLMs in various scientific fields across diverse modalities, it becomes necessary to comprehensively review related work in this direction. However, existing scientific LLM surveys typically focus on either one or two fields (e.g., biomedicine Wang et al. (2023a); He et al. (2024b); Pei et al. (2024); Zhang et al. (2024d) and chemistry Xia et al. (2023); Pei et al. (2024); Zhang et al. (2024d)) or one modality (e.g., text Ho et al. (2024)) only. In fact, if we take a holistic view of the research landscape, we can observe similar and interrelated techniques used to develop LLMs for different fields and modalities.
@@
Figure 1 depicts three major types of scientific LLM pre-training strategies (i.e., Columns 1 to 3), for each of which we give 4 examples (i.e., Types a to d). In Column 1, following BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), existing studies use masked language modeling (MLM) to pre-train encoder language models. Here, the input can be naturally sequential (e.g., papers in each field; protein, DNA, and RNA sequences in the FASTA format Lipman and Pearson (1985)) or artificially linearized (e.g., molecules in the SMILES format Weininger (1988); sequences of venue, author, and reference nodes in citation graphs). In Column 2, inspired by GPT Brown et al. (2020) and LLaMA Touvron et al. (2023a), previous studies adopt next token prediction to pre-train (encoder-)decoder language models, some of which further adopt instruction tuning and preference optimization Ouyang et al. (2022). Other than plain text input (e.g., question-answer pairs from knowledge bases or exams), we see more ways to sequentialize complex scientific data, such as flattening table cells and using particle coordinates to describe crystals. Even for images, there are studies in both mathematics Gao et al. (2023) and biomedicine Li et al. (2023a) that exploit a vision encoder to project an image onto several visual tokens and prepend them to text tokens as linearized LLM input. In Column 3, following DPR Karpukhin et al. (2020) and CLIP Radford et al. (2021), two encoders are pre-trained to map relevant data pairs closer in the latent space via contrastive learning. When both modalities are sequential (e.g., text-text or text-protein), the model is built upon two LLM encoders. When we prefer to keep the non-sequential nature of one modality (e.g., molecular graphs Edwards et al. (2021), chest X-rays Zhang et al. (2022), and aerial views Yan et al. (2024)), the corresponding graph or image encoder can be employed. To summarize, a cross-field, cross-modal survey will more accurately draw the connections between different scientific LLMs, demonstrate their commonalities, and potentially guide their future designs.
@@
Contributions. In this paper, motivated by the discussions above, we systematically survey over 260 scientific LLMs encompassing various fields (e.g., general science, mathematics, physics, chemistry, materials science, biology, medicine, and geoscience), modalities (e.g., language, graph, vision, table, molecule, protein, genome, and climate time series), and sizes (from $\sim$100M to $\sim$100B parameters). For each field/modality, we investigate commonly adopted pre-training datasets, model architectures, and evaluation tasks of scientific LLMs. Following our motivation, when we discuss model architectures in detail, we link them back to Figure 1 to build cross-field cross-modal connections. Moreover, we provide a structured summary of these scientific LLMs in Table A1-Table A6 (Appendix A). Furthermore, for different fields, we introduce how LLMs have been deployed to benefit science by augmenting different aspects and stages of the scientific discovery process, such as hypothesis generation, theorem proving, experiment design, drug discovery, and weather forecasting.
@@ 2.1 Language
General scientific LLMs are usually evaluated on common NLP tasks, such as named entity recognition (NER), relation extraction (RE) Luan et al. (2018), question answering (QA) Wang et al. (2024f), and classification Cohan et al. (2019).
@@ 2.3 Applications in Scientific Discovery
Performant scientific LLMs can work alongside researchers throughout the entire scientific discovery process. Leaving field-specific applications for later sections, here we underscore LLMs’ general usefulness in brainstorming and evaluation: Lahav et al. (2022) integrate LLMs into a search engine for the discovery of scientific challenges and directions; Wang et al. (2023e), Yang et al. (2024d), Baek et al. (2024), Gu and Krenn (2024), and Si et al. (2024) leverage LLMs to generate novel scientific ideas, directions, and hypotheses on the basis of prior literature and existing knowledge; Zhang et al. (2023h) rely on LLMs to find expert reviewers for each submission; Liu and Shah (2023), Liang et al. (2024c), and D’Arcy et al. (2024) explore the capacity of GPT-4 to provide useful feedback on research papers to facilitate automatic review generation; Liang et al. (2024b, a) also observe the increasing use of LLMs in writing scientific papers and conference peer reviews.
@@ 3.1 Language
The pre-training text corpora for mathematics LLMs can be categorized into two classes: (1) multiple-choice QA, the representative datasets of which include MathQA Amini et al. (2019), Ape210K Zhao et al. (2020), and Math23K Wang et al. (2017); as well as (2) generative QA, the representative datasets of which include GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021b), and MetaMathQA Yu et al. (2024c).
Similarly to general science LLMs, the backbone model of pioneering mathematics LLMs is BERT (Type 1.a, e.g., GenBERT Geva et al. (2020) and MathBERT Shen et al. (2021)), and these models are mostly trained via MLM. For GPT-based mathematics LLMs (Type 2.a, e.g., GSM8K-GPT Cobbe et al. (2021) and NaturalProver Welleck et al. (2022)), next token prediction and instruction tuning are major pre-training tasks to generate mathematical proofs and reasoning processes. The most recent models (Type 2.a, e.g., Rho-Math Lin et al. (2024b) and MAmmoTH2 Yue et al. (2024c)) are based on LLaMA and are trained to follow natural language instructions. However, when an enormous pre-training corpus is available (e.g., mathematical web pages and code), next token prediction is still favored as the mere pre-training task Azerbayev et al. (2024); Lin et al. (2024b) or the companion task Shao et al. (2024); Ying et al. (2024) to build base models.
QA and math world problems (MWP) have been the most common evaluation tasks for mathematics LLMs. In addition, quantitative reasoning contains more difficult problems, as the model has to provide a complete and self-contained solution without relying on external tools Shao et al. (2024); Lin et al. (2024b). We see a dominance of use from GSM8K and MATH for QA, and from MathQA and Math23K for MWP. For quantitative reasoning, MMLU-STEM Hendrycks et al. (2021a) and Big-Bench Hard Suzgun et al. (2023) are the most widely adopted.
@@ 3.2 Language + Vision
Geometry is one of the most important branches of mathematics, and it expresses the settings jointly in text and diagrams. As such, it is mandatory to involve the vision modality for geometry LLMs. The most commonly used pre-training datasets for geometry LLMs include Geometry3K Lu et al. (2021) and GeoQA Chen et al. (2021), both of which contain multiple-choice geometry problems.
@@
Geometry LLMs are evaluated through geometry problem solving, where the model is asked to select the correct answer given the diagram and its caption, the question, and answer options. Renowned evaluation datasets include Geometry3K Lu et al. (2021), GEOS Seo et al. (2015), and MathVista Lu et al. (2024).
@@ 3.3 Table
The challenge in tables is similar to that in diagrams, namely to obtain linearized table representations. In most cases, tables are squeezed into linear text sequences as part of the context and are prepended with the question text as the model input. As one of the first works in this line of research, TAPAS Herzig et al. (2020) (Type 1.a) adopts the MLM objective to predict the masked token in both textual and tabular contexts. Recent developments Li et al. (2024b); Zhang et al. (2024f) resemble the design of TableLlama Zhang et al. (2024e) (Type 2.b), with LLaMA-2 as the backbone and instruction tuning as the pre-training task.
@@ 3.4 Applications in Scientific Discovery
Mathematics LLMs have great potential to assist humans in offering potential solutions. For instance, AlphaGeometry Trinh et al. (2024) combines an LLM with a symbolic deduction engine, where the LLM generates useful constructs and the symbolic engine applies formal logic to find solutions. AlphaGeometry solves 25 out of 30 classical geometry problems adapted from the International Mathematical Olympiad. Sinha et al. (2024) extend AlphaGeometry by adding Wu’s method Chou (1988), further solving 27 out of 30, surpassing human gold medalists. FunSearch Romera-Paredes et al. (2024) integrates LLM with program search. One notable achievement of FunSearch is its ability to find a new solution to the cap set problem in combinatorial optimization. The solutions generated can be faster and more efficient than those devised by human experts. In Li et al. (2024a), LLMs iteratively propose and critique statistical models by leveraging in-context learning and chain-of-thought reasoning Wei et al. (2022b).
@@ 5.1 Language
LLM pre-training corpora in chemistry and materials science typically come from research papers and databases (e.g., Materials Project Jain et al. (2013)). Besides, recent works adopt domain-specific instruction tuning datasets (e.g., Mol-Instructions Fang et al. (2024a) and SMolInstruct Yu et al. (2024a)) derived from PubChem Kim et al. (2019), MoleculeNet Wu et al. (2018), etc.
@@ 5.2 Language + Graph
In some scenarios, molecular graphs appear simultaneously with text information, thus existing works have explored how to encode both effectively. The first type of such models adopts a GNN as the graph encoder and an LLM as the text encoder. The two modalities are connected through contrastive learning Liu et al. (2023d) (Type 3.c). For example, Text2Mol Edwards et al. (2021) uses GCN Kipf and Welling (2017) and SciBERT to encode a molecule and its corresponding natural language description, respectively, for text-to-molecule retrieval. The second type of such models utilizes an LLM to encode text and graphs simultaneously Zeng et al. (2022). Graphs can be either linearized to SMILES strings Edwards et al. (2022) (Type 2.c) or projected onto virtual tokens with graph encoders Zhao et al. (2023a); Liu et al. (2023e) (Type 2.d). For instance, 3D-MoLM Li et al. (2024c) uses a 3-dimensional molecular encoder to represent molecules as tokens and feeds them together with instructions into LLaMA-2 for molecule-to-text retrieval and molecule captioning.
@@ 5.4 Molecule
Different from subsection 5.2, this subsection introduces models dealing with molecules without associated text information. That being said, comparable approaches inspired by LLMs are utilized to develop molecular language models Flam-Shepherd et al. (2022). To be specific, most studies adopt SMILES or SELFIES Krenn et al. (2020) strings as the sequential representation of molecules. Similar to the trend in the “Language” modality, pioneering molecular LLMs focus on representation learning with bidirectional Transformer encoders (Type 1.c, e.g., SMILES-BERT Wang et al. (2019) and MoLFormer Ross et al. (2022)). For instance, ChemBERTa Chithrananda et al. (2020) adopts the architecture and pre-training strategy similar to those of RoBERTa Liu et al. (2019). These models exhibit extraordinary abilities in molecular understanding tasks such as molecular property prediction (e.g., toxicity classification Wu et al. (2018) and atomization energy regression Ramakrishnan et al. (2014)) as well as virtual screening Riniker and Landrum (2013). Later works explore the idea of representing molecules in an autoregressive fashion (Type 2.c, e.g., BARTSmiles Chilingaryan et al. (2024) and ChemGPT Frey et al. (2023)). For instance, T5Chem Lu and Zhang (2022) adopts the T5 backbone and a sequence-to-sequence pre-training objective. These models are evaluated in generative tasks that include molecule generation Gaulton et al. (2017), reaction prediction, and retrosynthesis. Besides linearizing molecules, there are studies modifying the Transformer architecture to admit molecular graphs, such as MAT Maziarka et al. (2020) and R-MAT Maziarka et al. (2024).
@@ 6.1 Language
The watershed moment in the evolution biomedical LLMs is still the emergence of billion-parameter architectures and instruction tuning. Before that, a wide variety of moderate-sized backbones are explored, including both encoder-based (Type 1.a, e.g., BioBERT Lee et al. (2020), Bio-ELECTRA Ozyurt (2020), BioRoBERTa Lewis et al. (2020b), BioALBERT Naseem et al. (2022), and Clinical-Longformer Li et al. (2022a)) and (encoder-)decoder-based ones (Type 2.a, e.g., SciFive Phan et al. (2021), BioBART Yuan et al. (2022a), and BioGPT Luo et al. (2022)). Evaluation tasks for these models range from biomedical NER, RE, sentence similarity estimation, document classification, and QA (i.e., the BLURB benchmark Gu et al. (2021)) to natural language inference (NLI) Romanov and Shivade (2018) and entity linking Doğan et al. (2014). After the watershed, the trend becomes instruction-tuning billion-parameter LLMs (Type 2.a, e.g., Med-PaLM Singhal et al. (2023a), MedAlpaca Han et al. (2023), and BioMistral Labrak et al. (2024)). Accordingly, evaluation tasks now include single-round QA Jin et al. (2021); Pal et al. (2022) and multi-round dialogue Wang et al. (2024g). Meanwhile, there are studies proposing a Bi-Encoder architecture (Type 3.a, e.g., Jin et al. (2023c) and Xu et al. (2024)) that specifically targets biomedical retrieval tasks, the benchmarks of which are NFCorpus Boteva et al. (2016), TREC-COVID Voorhees et al. (2021), etc.